If all the gas in a neon tank is used to fill 2500 balloons with a volume of 2.5L.. 8 what is the mass of neon in the tank?
Answer #1

The liters of all balloons are calculated:

V = 2500 balloons * 2.5 L / balloon = 6250 L

The mass is calculated by density:

m = V * d = 6250 L * 0.899 g / L = 5618.75 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.